# REST: Diffusion-based Real-time End-to-end Streaming Talking Head Generation via ID-Context Caching and Asynchronous Streaming Distillation

### v2

@@ 1 Introduction
Audio-driven Talking Head Generation (THG) aims to generate realistic talking head videos conditioned on input speech and reference images (Chen et al., 2020), demonstrating significant research and application value in domains such as remote education (Gowda et al., 2023), virtual reality (Zhen et al., 2023), and film production (Song et al., 2023). The lip-sync accuracy and naturalness of the generated videos serve as fundamental benchmarks for model performance (Song et al., 2023). Furthermore, due to growing demands of real-time applications like human-computer interaction, the inference speed of THG models is now garnering increasing research attention (Zhen et al., 2023).
@@
  • We propose REST, a novel framework for real-time, end-to-end, streaming talking head generation that effectively bridges the gap between diffusion and AR paradigms. To the best of our knowledge, REST is the first end-to-end method to achieve real-time, streaming audio-driven talking head diffusion on a single GPU.
@@ 2.2 Fast Talking Head Generation
Figure 2: The overall framework of REST. During training, the non-streaming teacher model is first pre-trained by asynchronous chunk-wise scheduler to provide a solid performance upper bound, shown in ①. Then the streaming student model with ID-Context Cache is trained under the guidance of the streaming teacher by ASD, shown in ②. The principle of ASD is demonstrated in ③
@@ 3.2 Streaming Diffusion with ID-Context Cache
where ${\mathit{\boldsymbol{X}}}(0)\in\mathbb{R}^{H\times W\times F\times D_{\text{v}}}$ denotes the video sequence, and ${\mathit{\boldsymbol{Z}}}(0)\in\mathbb{R}^{h\times w\times f\times d_{\text{v}}}$ are the compressed video latents. $\mathcal{E}_{\text{V}}$ and $\mathcal{D}_{\text{V}}$ denotes the encoder and decoder of VAE.
@@
Then the corresponding compact speech latent embedding $\mathit{\boldsymbol{E}\in\mathbb{R}^{f\times h_{w}\times d_{\text{A}}}}$ are generated by the SpeechAE encoder from $\mathit{\boldsymbol{S}}$, and reconstructed to $\hat{\mathit{\boldsymbol{S}}}$ through the decoder:
@@
where $H_{w}$ and $h_{w}$ indicates the window sizes, $D_{\text{A}}$ and $d_{\text{A}}$ denotes the hidden dims. Detailed architecture information is provided in Sec. B.2 of the Appendix.
@@
The mathematical formulation of the ID-Context Cache is defined as follows. For the current chunk, the Query, Key, and Value at each layer are computed as formulated below:
@@
where $i$ denotes the $i$-th chunk and $j$ denotes the $j$-th block of A2V-DiT. Then, the causal chain is reconstructed by concatenating the static ID anchor with the dynamically updated contextual cache from the preceding chunk, following the proposed ID-Sink and Context-Cache principles:
@@
Regarding cross-modal audio conditioning, the speech latent code $\boldsymbol{E}=[\boldsymbol{e}_{0}\mathbin{\|}\boldsymbol{e}_{1}\mathbin{\|}\cdots\mathbin{\|}\boldsymbol{e}_{k}]$ ($k$ denotes the number of chunks), which is temporally aligned with the video chunks, serves as the driving condition. By enforcing a frame-level spatial cross-attention with a temporal receptive field of chunk length, we achieve precise streaming audio-visual alignment within the compact latent space, formulated as:
$\mathit{\boldsymbol{H}}_{ij}^{\text{A}}=\mathit{\boldsymbol{H}}_{ij}+\text{CrossAttn}\left(\mathit{\boldsymbol{H}}_{ij},{\mathit{\boldsymbol{e}}_{j}}\right)$ (13)
where $\mathit{\boldsymbol{H}}_{ij},\mathit{\boldsymbol{H}}_{ij}^{\text{A}}\in\mathbb{R}^{h\times w\times f\times d}$ denotes the hidden states before and after frame-level spatial cross-attention of $i\!-\!\text{th}$ attention block. Our proposed design enables the semi-AR streaming generation of video latent chunks that are synchronized with the corresponding speech conditions. Additional analysis of the ID-Context Cache is provided in Sec. B.1.
@@ 3.3 Training with Asynchronous Streaming Distillation
$\boldsymbol{v}^{\mathcal{T}}$ $=\boldsymbol{\epsilon}-{\mathit{\boldsymbol{Z}^{\mathcal{T}}}}(0)$ (17)
@@
$\boldsymbol{v}^{\mathcal{S}}$ $=\boldsymbol{\epsilon}-\boldsymbol{Z}(0)$ (19)
@@
Teacher-Student Learning with ASD. There exists a fundamental information asymmetry between the teacher and student. The non-streaming teacher model leverages global attention to achieve superior consistency, while the streaming student is limited by local causal attention under streaming constraints. To address this fundamental limitation, we propose ASD between the non-streaming teacher and the streaming student. From an information-theoretic perspective, we aim to maximize the average mutual information (MI) between the generative distributions of student and teacher, thereby introducing global consistency constraints under streaming conditions. This is achieved by maximizing the MI variational lower bound via frame-wise contrastive learning, specifically by minimizing the InfoNCE loss (Oord et al., 2018) of the output velocity flows ${\mathit{\boldsymbol{v}^{\mathcal{T}}}}=\{{\mathit{\boldsymbol{v}}}^{\mathcal{T}}_{i}\}_{i=0}^{f}$ and ${\mathit{\boldsymbol{v}^{\mathcal{S}}}}=\{{\mathit{\boldsymbol{v}}}^{\mathcal{S}}_{i}\}_{i=0}^{f}$ (proof in Appendix Sec. C.1). The contrastive objective can be formulated as follows:
@@
Analysis of the loss complementarity is detailed in Sec. C.3. ASD serves to transfer global contextual knowledge from a non-streaming teacher to the streaming student, improving the stability and consistency of streaming generation.
@@ 4.1 Experimental Setup
Evaluation Metrics. Generation performance is assessed using several metrics. For visual quality, we employ the Fréchet Inception Distance (FID) (Seitzer, 2020) for image-level realism between synthesized videos and reference images and the Fréchet Video Distance (FVD) (Unterthiner et al., 2019) for frame-level realism between synthesized and ground-truth videos; lower values indicate better performance for both metrics. Lip synchronization is measured with SyncNet (Chung and Zisserman, 2017), where a higher Synchronization Confidence (Sync-C) and a lower Synchronization Distance (Sync-D) indicate superior alignment with speech input. We further use the Expression-FID (E-FID) metric from EMO (Tian et al., 2024) to measure the expression divergence between synthesized and ground-truth videos, with lower values indicating more faithful reproduction of expressions. Finally, we evaluate the efficiency of the diffusion backbone of each model by measuring the average runtime of the backbone per video (Runtime).
Baselines. We benchmark our method against several SOTA open-source methods, including end-to-end diffusion methods such as Sonic (Ji et al., 2025), EchoMimic (Chen et al., 2025), Hallo (Xu et al., 2024), FantasyTalking (Wang et al., 2025c) and AniPortrait (Wei et al., 2024), as well as motion-space diffusion methods like AniTalker (Liu et al., 2024) and Ditto (Li et al., 2025). All comparisons are conducted on the same device using identical test data with the same length of 4.84s (121 frames) to ensure fair evaluation.
@@ 4.2 Overall Comparison
As shown in Tab. 1, motion-space diffusion approaches such as Ditto and AniTalker generally demonstrate lower latency compared to end-to-end baselines. In contrast, our proposed end-to-end solution not only achieves a lower latency superior to all existing methods, but also pioneers streaming inference capabilities, which remains unattainable by existing alternatives. Beyond this efficiency, our approach achieves highly competitive performance, achieving either the best or second-best results across all metrics on both datasets, representing a significant breakthrough in real-time streaming talking head video generation.
@@ 4.3 Ablation Study
Ablation Study on ASD. To validate the effectiveness of our proposed ASD strategy and the contribution of each loss terms, we conduct a quantitative ablation study using the following configurations:
@@
We report the FID, FVD, and SyncNet scores of each configuration on the HDTF dataset, as shown in Tab 3. As indicated by the results, removing the Smooth loss while retaining the Contrastive loss for teacher-student distillation leads to a deterioration in the FVD metric, whereas FID and Sync-C scores show only marginal degradation. This indicates that the Smooth loss plays a crucial role in enhancing temporal coherence and preserving the overall semantic consistency of the generated videos. Conversely, removing the Contrastive loss and retaining only the Smooth loss results in noticeable deterioration in both Sync-C and FID metrics, with minimal impact on FVD. This demonstrates that the Contrastive loss primarily contributes to improving the overall visual quality and audio-visual alignment of the student model during streaming generation. Furthermore, the complete removal of the ASD strategy shows a substantial performance drop across all three metrics compared to the Full ASD configuration. These results validate the significant efficacy of our non-streaming to streaming distillation strategy in improving both video quality and temporal consistency of semi-autoregressive streaming generation.
@@ 4.4 Case Study
To provide a more intuitive comparison of the generation quality among SOTA THG methods, we present a visual analysis of a representative test case. Conditioned on the identical reference image and driving audio, we utilize each model to generate talking head videos of equal length of 121 frames. Key frames at identical timestamps are sampled for visual comparison, as illustrated in Fig. 4, with the inference runtime of backbone network for each model provided.
@@
Visual analysis of the representative case reveals distinct performance characteristics among THG methods. Motion-space diffusion methods such as AniTalker and Ditto achieve low runtime latency, yet exhibit poor lip-sync, with visible mismatches between generated and ground truth frames at corresponding timestamps (indicated by red arrows), underscoring the limitations of their two-stage training pipelines. End-to-end models, including Sonic and EchoMimic, show improved lip alignment, though occasional inaccuracies persist. Meanwhile, Hallo suffers from noticeable artifacts in later frames, attributable to its inferior generative quality. In contrast, our model operates with minimal inference time while maintaining accurate lip synchronization under real-time streaming constraints. The generated lip motions closely align with the ground truth, and the outputs exhibit natural head movements and facial expressions. Together, these results demonstrate that our approach achieves an optimal balance between efficiency and perceptual quality, confirming the effectiveness of the proposed method.

### v3

@@ 1 Introduction
Audio-driven Talking Head Generation (THG) aims to generate realistic talking head videos conditioned on input speech and reference images (Chen et al., 2020), demonstrating significant research and application value in domains such as remote education (Gowda et al., 2023), virtual reality (Zhen et al., 2023) and film production (Song et al., 2023). The lip-sync accuracy and naturalness of the generated videos serve as fundamental benchmarks for model performance (Song et al., 2023). Furthermore, due to the growing demands of real-time applications like human-computer interaction, the inference speed of THG models is garnering increasing research attention (Zhen et al., 2023).
@@
  • We propose REST, a novel framework for real-time, end-to-end, streaming talking head generation that effectively bridges the gap between diffusion and AR paradigms. To the best of our knowledge, REST is the first end-to-end method to achieve real-time streaming audio-driven talking head diffusion on a single GPU.
@@ 2.2 Fast Talking Head Generation
Figure 2: The overall framework of REST. During training, the non-streaming teacher model is first pre-trained with an asynchronous chunk-wise scheduler to provide a solid performance upper bound, shown in ①. Then the streaming student model with ID-Context Cache is trained under the guidance of the streaming teacher by ASD, as shown in ②. The principle of ASD is demonstrated in ③
@@ 3.2 Streaming Diffusion with ID-Context Cache
where ${\mathit{\boldsymbol{X}}}(0)\in\mathbb{R}^{H\times W\times F\times D_{\text{v}}}$ denotes the video sequence, and ${\mathit{\boldsymbol{Z}}}(0)\in\mathbb{R}^{h\times w\times f\times d_{\text{v}}}$ are the compressed video latents. $\mathcal{E}_{\text{V}}$ and $\mathcal{D}_{\text{V}}$ denote the encoder and decoder of VAE.
@@
Then the corresponding compact speech latent embedding $\mathit{\boldsymbol{E}\in\mathbb{R}^{f\times h_{w}\times d_{\text{A}}}}$ is generated by the SpeechAE encoder from $\mathit{\boldsymbol{S}}$, and reconstructed to $\hat{\mathit{\boldsymbol{S}}}$ through the decoder:
@@
where $H_{w}$ and $h_{w}$ indicate the window sizes, $D_{\text{A}}$ and $d_{\text{A}}$ denote the hidden dimensions. Detailed architecture information is provided in Sec. B.2 of the Appendix.
@@
The mathematical formulation of the ID-Context Cache is defined as follows. For the current chunk, the Query, Key, and Value at each block are computed as formulated below:
@@
where $i$ denotes the $i$-th chunk and $j$ denotes the $j$-th block of A2V-DiT. The causal chain is then reconstructed by concatenating the static ID anchor with the dynamically updated contextual cache from the preceding chunk, following the proposed ID-Sink and Context-Cache principles:
@@
Regarding cross-modal audio conditioning, the speech latent code $\boldsymbol{E}=[\boldsymbol{E}_{1}\mathbin{\|}\cdots\mathbin{\|}\boldsymbol{E}_{k}]$ ($k$ denotes the number of chunks), which is temporally aligned with the video chunks, serves as the driving condition. By enforcing frame-level spatial cross-attention with a temporal receptive field of chunk length $l$, we achieve precise streaming audio-visual alignment within the compact latent space, formulated as:
$\mathit{\boldsymbol{H}}_{ij}^{\text{A}}=\mathit{\boldsymbol{H}}_{ij}+\text{CrossAttn}\left(\mathit{\boldsymbol{H}}_{ij},{\mathit{\boldsymbol{E}}_{i}}\right)$ (13)
where $\mathit{\boldsymbol{H}}_{ij},\mathit{\boldsymbol{H}}_{ij}^{\text{A}}\in\mathbb{R}^{h\times w\times f\times d}$ denotes the hidden states before and after frame-level spatial cross-attention of $j\!-\!\text{th}$ attention block. Our proposed design enables the semi-AR streaming generation of video latent chunks that are synchronized with the corresponding speech conditions. Additional analysis of the ID-Context Cache is provided in Sec. B.1.
@@ 3.3 Training with Asynchronous Streaming Distillation
$\boldsymbol{v}$ $=\boldsymbol{\epsilon}-{\mathit{\boldsymbol{Z}}}(0)$ (17)
@@
$\boldsymbol{v}$ $=\boldsymbol{\epsilon}-\boldsymbol{Z}(0)$ (19)
@@
Teacher-Student Learning with ASD. There exists a fundamental information asymmetry between the teacher and the student. The non-streaming teacher model leverages global attention to achieve superior consistency, while the streaming student is limited by local causal attention under streaming constraints. To address this fundamental limitation, we propose ASD between the non-streaming teacher and the streaming student. From an information-theoretic perspective, we aim to maximize the average mutual information (MI) between the generative distributions of student and teacher, thereby introducing global consistency constraints under streaming conditions. This is achieved by maximizing the MI variational lower bound via frame-wise contrastive learning, specifically by minimizing the InfoNCE loss (Oord et al., 2018) of the output velocity flows ${\mathit{\boldsymbol{v}^{\mathcal{T}}}}=\{{\mathit{\boldsymbol{v}}}^{\mathcal{T}}_{i}\}_{i=0}^{f}$ and ${\mathit{\boldsymbol{v}^{\mathcal{S}}}}=\{{\mathit{\boldsymbol{v}}}^{\mathcal{S}}_{i}\}_{i=0}^{f}$ (proof in Appendix Sec. C.1). The contrastive objective can be formulated as follows:
@@
The analysis of loss complementarity is detailed in Sec. C.3. ASD serves to transfer global contextual knowledge from a non-streaming teacher to the streaming student, improving the stability and consistency of streaming generation.
@@ 4.1 Experimental Setup
Evaluation Metrics. Generation performance is assessed using several metrics. For visual quality, we employ the Fréchet Inception Distance (FID) (Seitzer, 2020) for image-level realism between synthesized videos and reference images and the Fréchet Video Distance (FVD) (Unterthiner et al., 2019) for frame-level realism between synthesized and ground-truth videos; lower values indicate better performance for both metrics. Lip synchronization is measured with SyncNet (Chung and Zisserman, 2017), where a higher Synchronization Confidence (Sync-C) and a lower Synchronization Distance (Sync-D) indicate superior alignment with speech input. We further use the Expression-FID (E-FID) metric from EMO (Tian et al., 2024) to measure the expression divergence between synthesized and ground-truth videos, with lower values indicating a more faithful reproduction of expressions. Finally, we evaluate the efficiency of the diffusion backbone of each model by measuring the average runtime of the backbone per video (Runtime).
Baselines. We benchmark our method against several open-source SOTA methods, including end-to-end diffusion methods such as Sonic (Ji et al., 2025), EchoMimic (Chen et al., 2025), Hallo (Xu et al., 2024), FantasyTalking (Wang et al., 2025c) and AniPortrait (Wei et al., 2024), as well as motion-space diffusion methods like AniTalker (Liu et al., 2024) and Ditto (Li et al., 2025). All comparisons are conducted on the same device using identical test data with the same length of 4.84s (121 frames) to ensure fair evaluation.
@@ 4.2 Overall Comparison
As shown in Tab. 1, motion-space diffusion approaches such as Ditto and AniTalker generally demonstrate lower latency compared to end-to-end baselines. In contrast, our proposed end-to-end solution not only achieves a lower latency superior to all existing methods, but also pioneers streaming inference capabilities, which remain unattainable by existing alternatives. Beyond this efficiency, our approach achieves highly competitive performance, achieving either the best or second-best results across all metrics on both datasets, representing a significant breakthrough in real-time streaming talking head video generation.
@@ 4.3 Ablation Study
Ablation Study on ASD. To validate the effectiveness of our proposed ASD strategy and the contribution of each loss term, we conduct a quantitative ablation study using the following configurations:
@@
We report the FID, FVD, and SyncNet scores of each configuration on the HDTF dataset, as shown in Tab 3. As indicated by the results, removing the Smooth loss while retaining the Contrastive loss for teacher-student distillation leads to a deterioration in the FVD metric, whereas FID and Sync-C scores show only marginal degradation. This indicates that the Smooth loss plays a crucial role in enhancing temporal coherence and preserving the overall semantic consistency of the generated videos. Conversely, removing the Contrastive loss and retaining only the Smooth loss results in noticeable deterioration in both Sync-C and FID metrics, with minimal impact on FVD. This demonstrates that the Contrastive loss primarily contributes to improving the overall visual quality and audio-visual alignment of the student model during streaming generation. Furthermore, complete removal of the ASD strategy shows a substantial performance drop across all three metrics compared to the Full ASD configuration. These results validate the significant efficacy of our non-streaming to streaming distillation strategy in improving both video quality and temporal consistency of semi-autoregressive streaming generation.
@@ 4.4 Case Study
To provide a more intuitive comparison of the generation quality among SOTA THG methods, we present a visual analysis of a representative test case. Conditioned on the identical reference image and driving audio, we utilize each model to generate talking head videos of equal length of 121 frames. Key frames at identical timestamps are sampled for visual comparison, as illustrated in Fig. 4, with the inference runtime of the backbone network for each model provided.
@@
Visual analysis of the representative case reveals distinct performance characteristics among THG methods. Motion-space diffusion methods such as AniTalker and Ditto achieve low runtime latency, yet exhibit poor lip-sync, with visible mismatches between generated and ground truth frames at corresponding timestamps (indicated by red arrows), underscoring the limitations of their two-stage training pipelines. End-to-end models, including Sonic and EchoMimic, show improved lip alignment, though occasional inaccuracies persist. Meanwhile, Hallo suffers from noticeable artifacts in later frames, attributable to its inferior generative quality. In contrast, our model operates with minimal inference time while maintaining accurate lip synchronization under real-time streaming constraints. The generated lip motions are closely aligned with the ground truth, and the outputs exhibit natural head movements and facial expressions. Together, these results demonstrate that our approach achieves an optimal balance between efficiency and perceptual quality, confirming the effectiveness of the proposed method.